OpenStudy (fibonaccichick666):

ok stuck yet again. I need to find the general solution of the linear system: \[{\huge{\{}} \left(\begin{matrix}x_1'=-5x_1-6x_2 \\ x_2'=2x_1+2x_2\end{matrix}\right)\] @ganeshie8 @Marki I'm in need of a nice hint

OpenStudy (anonymous):

\[{\huge{\{}} \left(\begin{matrix}x_1'=-5x_1-6x_2 \\ x_2'=2x_1+2x_2\end{matrix}\right)\]

OpenStudy (fibonaccichick666):

So I can find A but the issue is I don't have x(0) so I'm a bit confused. Is there a way other than Laplace transforms?

ganeshie8 (ganeshie8):

@Marki

ganeshie8 (ganeshie8):

if i remember, we find eigen values and eigen vectors and cookup the general solution ?

OpenStudy (fibonaccichick666):

I mean I can find those no problem, but I really can't figure out the approach here.

OpenStudy (anonymous):

yes yes